Consider the circuit shown below. (Due to the nature of this problem, do not use rounded intermediate values in your calculatplease answer asap
Answer #1

و ۱۱ 46 {ar Is Y Iu M 32 Ban I c Iua 4/2=2A At polut B G - Iz Ig=0.0 In loop EB AFE +32 = 11 € in loop EBCDE: T+2 Iz = 4 € s